# thanks: Corresponding author

### v1

@@ III Results and Discussion
This is precisely what we observe in the CIPSI-driven CC($P$) and CC($P$;$Q$) calculations of the singlet–triplet gaps, especially the latter ones, reported in Table 3 and Figure 2 (e) and (f). Indeed, the CC($P$;$Q$) approach using $N_{\text{det(in)}}=100,000$, which relies on small CIPSI diagonalization spaces [small $N_{\text{det(out)}}$ values], whose dimensionalities are about 1% of the $14,483,876$ $A_{g}(D_{\rm 2h})$-symmetric $S_{z}=0$ and $14,339,992$ $B_{1g}(D_{\rm 2h})$-symmetric $S_{z}=1$ triply excited amplitudes involved in the parent full CCSDT computations, and which employs even smaller $P$ spaces having only 0.1–0.6% of all triples, reduces the 0.553, 0.813, 1.300, 2.282, 4.434, and 9.222 kcal/mol differences between the CR-CC(2,3) and CCSDT $\Delta E_{\text{S--T}}$ values at $\lambda=0$, 0.2, 0.4, 0.6, 0.8, and 1 to 0.253, 0.301, 0.485, 0.784, 0.398, and 2.227 kcal/mol, respectively. When $N_{\text{det(in)}}$ is increased to $250,000$, where the numbers of determinants included in the final CIPSI diagonalizations preceding the CC($P$) and CC($P$;$Q$) steps are still only $\sim$2–3of all $T_{3}$ amplitudes used in the CCSDT calculations for the ${}^{1}A_{g}(D_{\rm 2h})$ and ${}^{3}B_{1g}(D_{\rm 2h})$ states and where the resulting $P$ spaces contain only 0.5–0.9% of all triples, the errors in the CC($P$;$Q$) $\Delta E_{\text{S--T}}$ gaps relative to their CCSDT counterparts at the above values of $\lambda$ decrease even more, to 0.112, 0.108, 0.128, 0.128, 0.152, and 0.263 kcal/mol, respectively, bringing the CC($P$;$Q$) and CCSDT results to a virtually perfect agreement, while reducing a computational effort compared to the CCSDT runs by orders of magnitude. As in the previously discussed results for the lowest singlet and triplet states of cyclobutadiene, especially for the challenging, $A_{g}(D_{\rm 2h})$-symmetric, singlet ground state, the $\delta(\mbox{$P$;$Q$})$ corrections play a major role in the observed error reductions, substantially improving the CC($P$) $\Delta E_{\text{S--T}}$ values, but, unlike in the CC($P$) calculations of the ${}^{1}A_{g}(D_{\rm 2h})$ and ${}^{3}B_{1g}(D_{\rm 2h})$ potentials, the singlet–triplet gaps obtained with the uncorrected CC($P$) approach using relatively small CIPSI diagonalization spaces and tiny fractions of all triples in the associated $P$ spaces can be quite accurate in their own right, reproducing the CCSDT values of $\Delta E_{\text{S--T}}$ across the entire $\lambda=0$–1 region to within $\sim$1–2 kcal/mol when $N_{\text{det(in)}}\gtrsim 250,000$. Clearly, this is a lot better than the 5.261, 8.589, and 15.120 kcal/mol errors relative to CCSDT obtained at $\lambda=0.6$, 0.8, and 1, respectively, with CCSD, demonstrating that the relaxation of $T_{1}$ and $T_{2}$ amplitudes in the presence of the leading $T_{3}$ contributions in the CIPSI-driven CC($P$) computations using sufficiently large $N_{\text{det(in)}}$ values results in a more balanced description of the many-electron correlation effects in the ${}^{1}A_{g}(D_{\rm 2h})$ and ${}^{3}B_{1g}(D_{\rm 2h})$ states, especially when $\lambda$ approaches 1, although, by inspecting Tables 13, we can also see that much of the improvement in the CCSD $\Delta E_{\text{S--T}}$ data offered by the CC($P$) approach originates from error cancellations between the ${}^{1}A_{g}(D_{\rm 2h})$ and ${}^{3}B_{1g}(D_{\rm 2h})$ CC($P$) energies. In the case of the CIPSI-driven CC($P$;$Q$) computations, we do not have to count on error cancellations, since both the total electronic energies of the lowest singlet and triplet states of cyclobutadiene at various values of $\lambda$ and the gaps between them rapidly converge toward their CCSDT parents with $N_{\text{det(in)}}$. The $\delta(\mbox{$P$;$Q$})$ corrections, in addition to being highly effective in improving the CC($P$) energies of the ${}^{1}A_{g}(D_{\rm 2h})$ and ${}^{3}B_{1g}(D_{\rm 2h})$ states and the associated $\Delta E_{\text{S--T}}$ values, are also very helpful in curing the nonsystematic error patterns in the singlet–triplet gaps observed in the CC($P$) calculations in Table 3 as $N_{\text{det(in)}}$ increases, where the differences between the CC($P$) and CCSDT values of $\Delta E_{\text{S--T}}$ go up and down or oscillate. This does not happen in the CC($P$;$Q$) calculations, where the resulting singlet–triplet gaps approach their CCSDT parents systematically and very fast as $N_{\text{det(in)}}$ is made larger, independent of $\lambda$. This is yet another demonstration of the ability of the CIPSI-driven CC($P$;$Q$) approach to provide a highly accurate and well-balanced description of the many-electron correlation effects characterizing the lowest-energy singlet and triplet states of cyclobutadiene along its automerization coordinate, which the conventional CC methods, such as CCSD, CR-CC(2,3), and other noniterative triples corrections to CCSD, cannot provide as one approaches the barrier region.
@@
We conclude this section by noticing that the results reported in Tables 13 also show that the convergence of the CIPSI-driven CC($P$) and CC($P$;$Q$) energies of the lowest ${}^{1}A_{g}(D_{\rm 2h})$ and ${}^{3}B_{1g}(D_{\rm 2h})$ states of cyclobutadiene and the gap between them toward their respective CCSDT parents with $N_{\text{det(in)}}$ or $N_{\text{det(out)}}$ is faster than that characterizing the associated variational and perturbatively corrected CIPSI energetics toward the extrapolated $E_{\text{var}}+\Delta E_{\text{r}}^{(2)}$ values. This observation is consistent with our initial study announcing the CIPSI-based CC($P$) and CC($P$;$Q$) methodologies Gururangan et al. (2021) and the fact that the CC($P$;$Q$) calculations are capable of accurately approximating the parent CCSDT energetics out of the unconverged CIPSI runs using relatively small Hamiltonian diagonalization spaces, even when $T_{3}$ correlations are large and nonperturbative and electronic quasi-degeneracies become substantial, as is the case in the barrier region of the ground-state ${}^{1}A_{g}(D_{2\text{h}})$ potential. While this may not be a general remark, we also observe that perturbatively corrected $E_{\text{var}}+\Delta E^{(2)}$ and $E_{\text{var}}+\Delta E^{(2)}_{r}$ energies of the lowest singlet and triplet states of cyclobutadiene and the gap between them converge toward their extrapolated limits at a rate similar to that characterizing our uncorrected CC($P$) calculations toward CCSDT. The $\delta(\mbox{$P$;$Q$})$-corrected CC($P$;$Q$) energetics converge to CCSDT much faster. This might be yet another way of looking at the effectiveness of the $\delta(\mbox{$P$;$Q$})$ moment corrections in improving the underlying CC($P$) results. That being said, we should keep in mind that the algorithms used to obtain the CCSDT and the perturbatively corrected and extrapolated CIPSI energies are fundamentally different procedures. Furthermore, and more importantly given the objectives of this study, where we are interested in exploring the CIPSI-driven CC($P$) and CC($P$;$Q$) methodologies, not the CIPSI approach itself, the CIPSI wave function growth in the calculations reported in Tables 13 was terminated long before our CIPSI runs were well converged, as we only needed information about the leading triply excited determinants and were not interested in saturating the triply excited manifolds of the relevant many-electron Hilbert spaces. Last but not least, to highlight the robustness of our CC($P$;$Q$) framework, all of the calculations reported in this work relied on the RHF and ROHF orbitals, i.e., we made no attempt to further optimize orbitals to make them consistent with correlated computations, which would improve CIPSI’s performance and which might also help our CC($P$) and CC($P$;$Q$) results using smaller $N_{\text{det(in)}}$ values, especially in the vicinity of the square TS geometry. We intend to look into potential benefits that might be offered by orbital optimizations in the CIPSI-driven CC($P$) and CC($P$;$Q$) calculations in a future study.
@@ IV Summary and Concluding Remarks
An accurate determination of singlet–triplet gaps in biradicals represents a formidable test for ab initio electronic structure methodologies, as it requires balancing strong nondynamical many-electron correlation effects, needed for a reliable description of the low-spin singlet states that have a manifestly multiconfigurational nature, with the generally weaker, largely dynamical, correlations characterizing the high-spin triplet states. Although high-level CC methods with a full treatment of higher–than–two-body clusters, such as CCSDT or CCSDTQ, are often powerful enough to capture the dynamical and nondynamical correlation effects relevant in such studies, their applications are hindered by the demanding computational steps and memory requirements, which are prohibitively expensive when larger many-electron systems are examined. One of the promising ideas aimed at addressing this situation within the single-reference CC framework is the CC($P$;$Q$) formalism, in which one solves the CC amplitude equations in a suitably defined subspace of the many-electron Hilbert space, referred to as the $P$ space, and improves the resulting CC($P$) energies using the a posteriori moment corrections, designated as $\delta(P;Q)$, calculated with the help of the complementary $Q$ space.Shen and Piecuch (2012, 2012, 2012); Bauman et al. (2017); Magoulas et al. (2018); Yuwono et al. (2019); Deustua et al. (2017); Yuwono et al. (2020); Deustua et al. (2021); Chakraborty et al. (2022); Gururangan et al. (2021); Gururangan and Piecuch (2023) Among the most attractive features of the CC($P$;$Q$) methodology is its flexibility, so that in addition to conventional choices of the $P$ and $Q$ spaces using truncations based on excitation ranks, which in the past resulted in the development of the biorthogonal CR-CC methods, such as the CR-CC(2,3) triples correction to CCSD, Piecuch and Włoch (2005); Piecuch et al. (2006); Włoch et al. (2006, 2007) one can consider various unconventional ways of setting up these spaces that can improve the CR-CC(2,3), CCSD(T), and similar energetics for systems with substantial electronic quasi-degeneracies by relaxing the $T_{1}$ and $T_{2}$ components of the cluster operator $T$ in the presence of their higher-rank $T_{n}$ counterparts with $n>2$, such as $T_{3}$, which become large, nonperturbative, and strongly coupled to $T_{1}$ and $T_{2}$ in such situations. This can be done without major increases in the computational effort by incorporating the leading higher–than–doubly excited determinants in the $P$ spaces and using corrections $\delta(P;Q)$ to capture the remaining correlations of interest.

### v2

@@ III Results and Discussion
This is precisely what we observe in the CIPSI-driven CC($P$) and CC($P$;$Q$) calculations of the singlet–triplet gaps, especially the latter ones, reported in Table 3, Figure 4, and Figure 5 (e) and (f). Indeed, the CC($P$;$Q$) approach using the cc-pVDZ basis and $N_{\text{det(in)}}=100,000$, which relies on small CIPSI diagonalization spaces [small $N_{\text{det(out)}}$ values], whose dimensionalities are about 1% of the $14,483,876$ $A_{g}(D_{\rm 2h})$-symmetric $S_{z}=0$ and $14,339,992$ $B_{1g}(D_{\rm 2h})$-symmetric $S_{z}=1$ triply excited amplitudes involved in the parent full CCSDT computations, and which employs even smaller $P$ spaces having only 0.1–0.6% of all triples, reduces the 0.553, 0.813, 1.300, 2.282, 4.434, and 9.222 kcal/mol differences between the CR-CC(2,3)/cc-pVDZ and CCSDT/cc-pVDZ $\Delta E_{\text{S--T}}$ values at $\lambda=0$, 0.2, 0.4, 0.6, 0.8, and 1 to 0.253, 0.301, 0.485, 0.784, 0.398, and 2.227 kcal/mol, respectively. When $N_{\text{det(in)}}$ is increased to $250,000$, where the numbers of determinants included in the final CIPSI diagonalizations preceding the CC($P$) and CC($P$;$Q$) steps are still only $\sim$2–3% of all $T_{3}$ amplitudes used in the target CCSDT/cc-pVDZ calculations for the $^{1}A_{g}(D_{\rm 2h})$ and $^{3}B_{1g}(D_{\rm 2h})$ states, and where the resulting $P$ spaces contain only 0.5–0.9% of all triples, the errors in the CC($P$;$Q$) $\Delta E_{\text{S--T}}$ gaps relative to their CCSDT counterparts obtained at the above values of $\lambda$ with the cc-pVDZ basis set decrease even more, to 0.112, 0.108, 0.128, 0.128, 0.152, and 0.263 kcal/mol, respectively, bringing the CC($P$;$Q$) and CCSDT results to a virtually perfect agreement, while reducing a computational effort compared to the CCSDT runs by orders of magnitude. As in the previously discussed results for the lowest singlet and triplet states of cyclobutadiene, especially for the challenging, $A_{g}(D_{\rm 2h})$-symmetric, singlet ground state, the $\delta(\mbox{$P$;$Q$})$ corrections play a major role in the observed error reductions, substantially improving the CC($P$) $\Delta E_{\text{S--T}}$ values, but, unlike in the CC($P$) calculations of the $^{1}A_{g}(D_{\rm 2h})$ and $^{3}B_{1g}(D_{\rm 2h})$ potentials, the singlet–triplet gaps obtained with the uncorrected CC($P$) approach using relatively small CIPSI diagonalization spaces and tiny fractions of all triples in the associated $P$ spaces can be quite accurate in their own right, reproducing the CCSDT values of $\Delta E_{\text{S--T}}$ across the entire $\lambda=0$–1 region to within $\sim$1–2 kcal/mol when $N_{\text{det(in)}}\gtrsim 250,000$ and the cc-pVDZ basis set is employed. Clearly, this is a lot better than the 5.261, 8.589, and 15.120 kcal/mol errors relative to CCSDT obtained at $\lambda=0.6$, 0.8, and 1, respectively, with CCSD using the same basis, demonstrating that the relaxation of $T_{1}$ and $T_{2}$ amplitudes in the presence of the leading $T_{3}$ contributions in the CIPSI-driven CC($P$) computations using sufficiently large $N_{\text{det(in)}}$ values results in a more balanced description of the many-electron correlation effects in the $^{1}A_{g}(D_{\rm 2h})$ and $^{3}B_{1g}(D_{\rm 2h})$ states, especially when $\lambda$ approaches 1, although, by inspecting Tables 13 and Figures 24, we can also see that much of the improvement in the CCSD $\Delta E_{\text{S--T}}$ data offered by the CC($P$) approach originates from error cancellations between the $^{1}A_{g}(D_{\rm 2h})$ and $^{3}B_{1g}(D_{\rm 2h})$ CC($P$) energies. In the case of the CIPSI-driven CC($P$;$Q$) computations, we do not have to count on error cancellations, since both the total electronic energies of the lowest singlet and triplet states of cyclobutadiene at various values of $\lambda$ and the gaps between them rapidly converge toward their CCSDT parents with $N_{\text{det(in)}}$. The $\delta(\mbox{$P$;$Q$})$ corrections, in addition to being highly effective in improving the CC($P$) energies of the $^{1}A_{g}(D_{\rm 2h})$ and $^{3}B_{1g}(D_{\rm 2h})$ states and the associated $\Delta E_{\text{S--T}}$ values, are also very helpful in curing the nonsystematic error patterns in the singlet–triplet gaps observed in the CC($P$) calculations in Table 3 and Figure 4 as $N_{\text{det(in)}}$ increases, where the differences between the CC($P$) and CCSDT values of $\Delta E_{\text{S--T}}$ go up and down or oscillate. This does not happen in the CC($P$;$Q$) calculations, where the resulting singlet–triplet gaps approach their CCSDT parents systematically and very fast as $N_{\text{det(in)}}$ is made larger, independent of $\lambda$. This is yet another demonstration of the ability of the CIPSI-driven CC($P$;$Q$) approach to provide a highly accurate and well-balanced description of the many-electron correlation effects characterizing the lowest-energy singlet and triplet states of cyclobutadiene along its automerization coordinate, which the conventional CC methods, such as CCSD, CR-CC(2,3), and other noniterative triples corrections to CCSD, cannot provide as one approaches the barrier region.
Given our generally positive experiences with the active-orbital-based variant of the CC($P$;$Q$) methodology abbreviated as CC(t;3),Shen and Piecuch (2012, 2012, 2012); Bauman et al. (2017); Magoulas et al. (2018); Yuwono et al. (2019); Gururangan et al. (2025) which corrects the CCSDt energetics for those $T_{3}$ correlations that are not captured by the active-space CCSDt approach, it is interesting to compare the lowest $^{1}A_{g}(D_{\rm 2h})$ and $^{3}B_{1g}(D_{\rm 2h})$ potentials of cyclobutadiene and the gap between them obtained in the CIPSI-driven CC($P$) and CC($P$;$Q$) calculations with their CCSDt and CC(t;3) counterparts. This is done in Figure 6, which compares the PECs corresponding to the lowest singlet and triplet states of cyclobutadiene, as described by the cc-pVDZ basis set, along the $D_{\rm 2h}$-symmetric automerization pathway defined by Eq. (2) resulting from the CIPSI-based CC($P$) and CC($P$;$Q$) calculations employing $N_{\text{det(in)}}=250,000$ with the analogous potentials generated with the active-orbital-based CCSDt and CC(t;3) methods and full CCSDT. The numerical data used to construct the CCSDt and CC(t;3) PECs shown in Figure 6 and to determine the gap between them can be found in Tables S1–S3 of the Supporting Information. We recall that in the language of the CC($P$) and CC($P$;$Q$) formalisms, the $P$ space adopted in the CCSDt iterations consists of all singly and doubly excited determinants and the subset of triply excited determinants that fit the formula $|\Phi_{ij{\bf K}}^{{\bf A}bc}\rangle$, where $i,j$ ($b,c$) designate the spinorbitals occupied (unoccupied) in the reference determinant $|\Phi\rangle$ and $\bf K$} ($\bf A$}) are the occupied (unoccupied) spinorbitals around the Fermi level belonging to the user-specified active set, Oliphant and Adamowicz (1992, 1993); Piecuch et al. (1993); Piecuch and Adamowicz (1994, 1995); Ghose et al. (1995); Adamowicz et al. (1998); Piecuch et al. (1999, 1999); Piecuch (2010) and the complementary $Q$ space needed to determine the CC(t;3) correction to CCSDt using Eq. (7) is spanned by the remaining triply excited determinants $|\Phi_{ijk}^{abc}\rangle$ outside the $|\Phi_{ij{\bf K}}^{{\bf A}bc}\rangle$ set. In the specific case of the CCSDt and CC(t;3) computations reported in Figure 6 and Tables S1–S3 of the Supporting Information, the active space defining the subsets of triply excited determinants included in the CCSDt calculations preceding the determination of the noniterative CC(t;3) corrections consisted of two orbitals of cyclobutadiene that correlate with the valence $e_{g}$ shell of the $D_{4\text{h}}$-symmetric TS ($\lambda=1$) structure, meaning the highest occupied and lowest unoccupied RHF orbitals for the lowest $^{1}A_{g}(D_{\rm 2h})$ state and the two singly occupied ROHF orbitals in the case of the lowest $^{3}B_{1g}(D_{\rm 2h})$ state. As explained in the Supporting Information (see footnotes ‘b’ in Tables S1 and S2), with these choices of active orbitals, the $P$ space used in the CCSDt computations for the lowest singlet state contained 1.5% of the $S_{z}=0$ triples of the $A_{g}(D_{\rm 2h})$ symmetry involved in the parent full CCSDT work, whereas that for the lowest triplet state contained 1.1% of all $S_{z}=1$ $B_{1g}(D_{\rm 2h})$-symmetric triples. These percentages should be compared to 0.5–0.6% and 0.6–0.9% of triples of the $S_{z}=0$ $A_{g}(D_{\rm 2h})$ and $S_{z}=1$ $B_{1g}(D_{\rm 2h})$ symmetries, respectively, captured by the CIPSI runs preceding the CC($P$) and CC($P$;$Q$) calculations with $N_{\text{det(in)}}=250,000$, which, as shown in Figure 6 (cf., also, Tables S1–S3 of the Supporting Information and Tables 13), produce the results that are in generally very good agreement with their CCSDt and CC(t;3) counterparts. One might argue that in spite of having a somewhat larger fraction of triply excited determinants in the $P$ spaces employed in the CCSDt calculations compared to the numbers of triples identified by the $N_{\text{det(in)}}=250,000$ CIPSI runs, the CC(t;3) results for the lowest-energy $^{3}B_{1g}(D_{\rm 2h})$ state reported in Table S2 of the Supporting Information are less accurate than those obtained with the CIPSI-based CC($P$;$Q$) approach using $N_{\text{det(in)}}=250,000$ (or even smaller $N_{\text{det(in)}}$ values) shown in Table 2. One might also argue that with the exception of the TS region, the usage of a larger fraction of triples in the $P$ spaces employed in the CCSDt computations for the lowest $^{1}A_{g}(D_{\rm 2h})$ state compared to the numbers of the triples identified by CIPSI using $N_{\text{det(in)}}=250,000$ does not translate into substantial improvements in the CIPSI-driven CC($P$;$Q$) results based on this $N_{\text{det(in)}}$ value by CC(t;3) (cf. Table S1 in the Supporting Information and Table 1). None of this, however, alters our conclusion regarding the generally good agreement between the CIPSI-driven CC($P$) and CC($P$;$Q$) calculations for the lowest singlet and triplet potentials of cyclobutadiene, as described by the cc-pVDZ basis, and the gap between them obtained with $N_{\text{det(in)}}=250,000$ and their CCSDt and CC(t;3) counterparts, and none of this is surprising. Indeed, our choice of active orbitals in the CCSDt and CC(t;3) calculations discussed here reflects on the multi-configurational character of the singlet ground state in the vicinity of the $\lambda=1$ TS region, but is not necessarily best for the lowest triplet state, which is dominated by dynamical correlations at all values of $\lambda$, or the singlet ground state as $\lambda\rightarrow 0$, where dynamical correlations dominate as well. Furthermore, unlike in the CCSDt case, the lists of triply excited determinants included in the $P$ spaces adopted in the CIPSI-driven CC($P$) calculations preceding the determination of the CC($P$;$Q$) corrections may vary with the nuclear geometry (adjusting to the wave function content as the nuclear geometry changes), so they tend to be more compact than those employed by CCSDt if the active-orbital-based CC(t;3) and CIPSI-based CC($P$;$Q$) computations become similarly accurate [cf. Ref. Gururangan et al., 2025 for the analogous comments regarding the adaptive CC($P$;$Q$) framework vs. CC(t;3)]. Having said all this, the similarity between the PECs characterizing the lowest singlet and triplet states of cyclobutadiene, as described by the cc-pVDZ basis, along its automerization pathway resulting from the CIPSI-driven CC($P$;$Q$) calculations using $N_{\text{det(in)}}=250,000$ and their counterparts obtained with the CC(t;3) approach using a chemically motivated active space consisting of two valence orbitals, combined with the observation that these two independent computations accurately approximate the parent CCSDT data, is reassuring. It demonstrates that while there may be some differences between the subsets of triply excited determinants identified with the help of active orbitals following the CCSDt recipe and those extracted from the CIPSI runs, these differences are relatively small if the terminal diagonalization space used by the CIPSI approach that drives the CC($P$) and CC($P$;$Q$) computations is sufficiently large and the active space used to set up the CCSDt calculations preceding the determination of the CC(t;3) corrections is reasonable. In other words, both CCSDt and CIPSI are capable of capturing the leading triples for inclusion in the $P$ spaces used by the CC($P$) and CC($P$;$Q$) calculations, but CIPSI allows us to do it in a more black-box fashion, without having to resort to user-defined active orbitals, which is certainly appealing. This remark is in line with one of our earlier studies, reported in Ref. Yuwono et al., 2020, where we compared the manifolds of triply excited determinants captured in the context of the semi-stochastic CC($P$;$Q$) considerations by CIQMC with those defined by the $|\Phi_{ij{\bf K}}^{{\bf A}bc}\rangle$ formula of CCSDt.
[…]
@@
We conclude this section by noticing that the results reported in Tables 13 for the cc-pVDZ basis and 46 for cc-pVTZ also show that the convergence of the CIPSI-driven CC($P$) and CC($P$;$Q$) energies of the lowest $^{1}A_{g}(D_{\rm 2h})$ and $^{3}B_{1g}(D_{\rm 2h})$ states of cyclobutadiene and the gap between them toward their respective CCSDT parents with $N_{\text{det(in)}}$ or $N_{\text{det(out)}}$ is faster than that characterizing the associated variational and perturbatively corrected CIPSI energetics toward the extrapolated $E_{\text{var}}+\Delta E_{\text{r}}^{(2)}$ values (this is particularly true for the larger cc-pVTZ basis set, where we would have to consider much larger $N_{\text{det(in)}}$ values and diagonalization spaces in CIPSI than those used in this study to obtain more accurate estimates of the extrapolated $E_{\text{var}}+\Delta E_{\text{r}}^{(2)}$ energies). This observation is consistent with our initial study announcing the CIPSI-based CC($P$) and CC($P$;$Q$) methodologies Gururangan et al. (2021) and the fact that the CC($P$;$Q$) calculations are capable of accurately approximating the parent CCSDT energetics out of the unconverged CIPSI runs using relatively small Hamiltonian diagonalization spaces, even when $T_{3}$ correlations are large and nonperturbative and electronic quasi-degeneracies become substantial, as is the case in the barrier region of the ground-state $^{1}A_{g}(D_{2\text{h}})$ potential. While this may not be a general remark, we also observe that perturbatively corrected $E_{\text{var}}+\Delta E^{(2)}$ and $E_{\text{var}}+\Delta E^{(2)}_{r}$ energies of the lowest singlet and triplet states of cyclobutadiene and the gap between them converge toward their extrapolated limits at a rate similar to that characterizing our uncorrected CC($P$) calculations toward CCSDT. The $\delta(\mbox{$P$;$Q$})$-corrected CC($P$;$Q$) energetics converge to CCSDT much faster. This might be yet another way of looking at the effectiveness of the $\delta(\mbox{$P$;$Q$})$ moment corrections in improving the underlying CC($P$) results. That being said, we should keep in mind that the algorithms used to obtain the CCSDT and the perturbatively corrected and extrapolated CIPSI energies are fundamentally different procedures. Furthermore, and more importantly given the objectives of this study, where we are interested in exploring the CIPSI-driven CC($P$) and CC($P$;$Q$) methodologies, not the CIPSI approach itself, the CIPSI wave function growth in the calculations reported in Tables 16 (especially in Tables 46 for the cc-pVTZ basis) was terminated long before our CIPSI runs were well converged, as we only needed information about the leading triply excited determinants and were not interested in saturating the triply excited manifolds of the relevant many-electron Hilbert spaces. Last but not least, to highlight the robustness of our CC($P$;$Q$) framework, all of the calculations reported in this work relied on the RHF and ROHF orbitals, i.e., we made no attempt to further optimize orbitals to make them consistent with correlated computations, which would improve CIPSI’s performance and which might also help our CC($P$) and CC($P$;$Q$) results using smaller $N_{\text{det(in)}}$ values, especially in the vicinity of the square TS geometry. We intend to look into potential benefits that might be offered by orbital optimizations in the CIPSI-driven CC($P$) and CC($P$;$Q$) calculations in a future study.
@@ IV Summary and Concluding Remarks
An accurate determination of singlet–triplet gaps in biradicals represents a formidable test for ab initio electronic structure methodologies, as it requires balancing strong nondynamical many-electron correlation effects, needed for a reliable description of the low-spin singlet states that have a manifestly multiconfigurational nature, with the generally weaker, largely dynamical, correlations characterizing the high-spin triplet states. Although high-level CC methods with a full treatment of higher–than–two-body clusters, such as CCSDT or CCSDTQ, are often powerful enough to capture the dynamical and nondynamical correlation effects relevant in such studies, their applications are hindered by the demanding computational steps and memory requirements, which are prohibitively expensive when larger many-electron systems are examined. One of the promising ideas aimed at addressing this situation within the single-reference CC framework is the CC($P$;$Q$) formalism, in which one solves the CC amplitude equations in a suitably defined subspace of the many-electron Hilbert space, referred to as the $P$ space, and improves the resulting CC($P$) energies using the a posteriori moment corrections, designated as $\delta(P;Q)$, calculated with the help of the complementary $Q$ space.Shen and Piecuch (2012, 2012, 2012); Bauman et al. (2017); Magoulas et al. (2018); Yuwono et al. (2019); Deustua et al. (2017); Yuwono et al. (2020); Deustua et al. (2021); Chakraborty et al. (2022); Gururangan et al. (2021); Gururangan and Piecuch (2023); Gururangan et al. (2025) Among the most attractive features of the CC($P$;$Q$) methodology is its flexibility, so that in addition to conventional choices of the $P$ and $Q$ spaces using truncations based on excitation ranks, which in the past resulted in the development of the biorthogonal CR-CC methods, such as the CR-CC(2,3) triples correction to CCSD, Piecuch and Włoch (2005); Piecuch et al. (2006); Włoch et al. (2006, 2007) one can consider various unconventional ways of setting up these spaces that can improve the CR-CC(2,3), CCSD(T), and similar energetics for systems with substantial electronic quasi-degeneracies by relaxing the $T_{1}$ and $T_{2}$ components of the cluster operator $T$ in the presence of their higher-rank $T_{n}$ counterparts with $n>2$, such as $T_{3}$, which become large, nonperturbative, and strongly coupled to $T_{1}$ and $T_{2}$ in such situations. This can be done without major increases in the computational effort by incorporating the leading higher–than–doubly excited determinants in the $P$ spaces and using corrections $\delta(P;Q)$ to capture the remaining correlations of interest.